### v1

# Topic Modeling for Short Texts with Large Language Models

Abstract
As conventional topic models rely on word co-occurrence to infer latent topics, topic modeling for short texts has been a long-standing challenge. Large Language Models (LLMs) can potentially overcome this challenge by contextually learning the semantics of words via pretraining. This paper studies two approaches, parallel prompting and sequential prompting, to use LLMs for topic modeling. Due to the input length limitations, LLMs cannot process many texts at once. By splitting the texts into smaller subsets and processing them parallelly or sequentially, an arbitrary number of texts can be handled by LLMs. Experimental results demonstrated that our methods can identify more coherent topics than existing ones while maintaining the diversity of the induced topics. Furthermore, we found that the inferred topics adequately covered the input texts, while hallucinated topics were hardly generated.
Topic Modeling for Short Texts with Large Language Models
Tomoki Doi The University of Tokyo doi-tomoki701@is.s.u-tokyo.ac.jp Masaru Isonuma The University of Edinburgh m.isonuma@ed.ac.uk Hitomi Yanaka The University of Tokyo hyanaka@is.s.u-tokyo.ac.jp
1 Introduction
Topic modeling is a classical task of discovering latent topics that best describe a set of documents Blei et al. (2003); Churchill and Singh (2022). Recently, while neural topic models work successfully on various kinds of long documents Miao et al. (2017); Srivastava and Sutton (2017); Dieng et al. (2020a), they can not handle short texts, such as social media posts and news headlines Li et al. (2016); Wu et al. (2022).
Large Language Models (LLMs), such as InstructGPT Ouyang et al. (2022) and GPT-4 OpenAI (2023), have shown impressive results on various tasks by providing task instructions in a zero-shot manner (Wang et al., 2023; Kocoń et al., 2023). Since conventional topic models infer the topics of words by relying on word co-occurrence, they perform worse on short texts. In contrast, as LLMs contextually learn the semantics of words by pretraining on massive text corpora, they could accurately infer the latent topics of short texts.
In order to use LLMs for topic modeling, we propose two approaches, parallel prompting and sequential prompting (Figure 1). Due to the input length limitations of LLMs, an input document set must be split into smaller subsets, which are processed individually. Parallel prompting concurrently infers the topics of each subset and merges them to represent the topics of the whole document set. Sequential prompting processes each subset successively, updating the topics in every iteration. We assess our approaches across texts from various domains using multiple evaluation metrics.
The contributions of this study are as follows:
1.
We propose parallel and sequential prompting methods for topic modeling using LLMs. Our methods can handle a large number of texts that cannot be processed in a single run due to the input length limitations of LLMs.
2.
We validate the performance of our methods by comparing them to existing models and show that ours can identify more coherent topics than existing models while maintaining the diversity of the induced topics.
3.
We assess the document coverage and factuality of the inferred topics, due to concerns that LLMs may focus on only parts of documents or generate hallucinated topics. Evaluation results indicate that our concerns are negligible.
2 Background
Topic modeling is a task of identifying latent topics as a set of topic words representing each topic from a collection of documents Blei et al. (2003). Topic modeling has conventionally been tackled with probabilistic models such as latent Dirichlet allocation (LDA, Blei et al., 2003). In recent years, however, neural models have become popular due to their high performance Srivastava and Sutton (2017); Dieng et al. (2020b); Grootendorst (2022).
It is known that topic modeling for short texts is difficult for current topic models due to data sparsity Li et al. (2016); Wu et al. (2022). TSCTM Wu et al. (2022) is a current state-of-the-art neural topic model for short texts. This model addresses data sparsity by learning representations of documents based on VQ-VAE van den Oord et al. (2017), contrastive learning, and incorporation of data augmentation into the learning.
BERTopic Grootendorst (2022) uses a pretrained encoder, Sentence-BERT Reimers and Gurevych (2019), to obtain clusters of documents and assigns topic words to each cluster by a class-based TF-IDF procedure. Another related study is Stammbach et al. (2023), which utilizes LLMs to evaluate topic quality automatically. However, our study is the first to explore how well LLMs perform topic modeling.
3 Topic Modeling with LLMs
We introduce two approaches to perform topic modeling with LLMs: parallel prompting and sequential prompting. For these approaches, we apply common preprocessing, which involves randomly splitting a document set into subsets with the same size, smaller than the context length of the LLMs.
Parallel Prompting
In the parallel prompting, LLMs identify topics for each subset in parallel by prompting the subset and the instruction of topic modeling. Then, the topics of each subset are merged by LLMs. We use two kinds of prompts shown in Table 1: (i) a Par prompt for parallel topic modeling for each subset, and (ii) a Par prompt for merging topics of each results.
Sequential Prompting
In the sequential prompting, LLMs identify topics for each subset sequentially, considering the topics previously identified for the previous subset. We use a Par for the first subset, then use a Seq prompt in Table 1 for the other subsets. This prompt contains topics identified in the prior subset and instructions to refer to them.
| ID | Prompt |
| Par | Write the results of simulating topic modeling |
| for the following documents: [DOCS]. | |
| Par | Write the results of merging the following |
| topic modeling results:[TOPICS],[TOPICS], … | |
| Seq | Write the results of simulating topic modeling |
| for the following documents: [DOCS], Make the | |
| most use of the following topics: [TOPICS]. |
| # of | Text | Vocabulary | |
| Dataset | Documents | Length | Size |
| Tweet | 2000 | 5.47 | 706 |
| GoogleNewsT | 11000 | 5.25 | 2376 |
| StackOverFlow | 19000 | 4.71 | 2544 |
| Model | Tweet | GoogleNewsT | StackOverFlow | |||||||||||
| Cv | TU | Cv | TU | Cv | TU | Cv | TU | Cv | TU | Cv | TU | |||
| LDA | 0.394 | 0.800 | 0.401 | 0.568 | 0.426 | 0.984 | 0.406 | 0.963 | 0.320 | 0.928 | 0.425 | 0.883 | ||
| LDAAug | 0.445 | 0.968 | 0.436 | 0.856 | 0.411 | 0.984 | 0.381 | 0.981 | 0.360 | 0.920 | 0.508 | 0.952 | ||
| TSCTM | 0.393 | 1.000 | 0.467 | 0.997 | 0.333 | 1.000 | 0.374 | 1.000 | 0.244 | 1.000 | 0.313 | 1.000 | ||
| TSCTMAug | 0.355 | 1.000 | 0.433 | 1.000 | 0.243 | 1.000 | 0.346 | 1.000 | 0.218 | 1.000 | 0.276 | 1.000 | ||
| BERTopic | 0.514 | 1.000 | 0.537 | 1.000 | 0.439 | 1.000 | 0.437 | 1.000 | 0.459 | 1.000 | 0.485 | 0.971 | ||
| BERTopicAug | 0.535 | 1.000 | 0.526 | 1.000 | 0.412 | 1.000 | 0.417 | 1.000 | 0.460 | 1.000 | 0.489 | 0.955 | ||
| GPT-3.5Seq | 0.552 | 0.960 | 0.515 | 0.920 | 0.562 | 0.984 | 0.489 | 0.948 | 0.441 | 0.896 | 0.517 | 0.775 | ||
| GPT-3.5Par | 0.476 | 0.992 | 0.532 | 0.900 | 0.571 | 0.960 | 0.535 | 0.913 | 0.312 | 0.864 | 0.496 | 0.913 | ||
| GPT-4Seq | 0.577 | 0.992 | 0.551 | 0.976 | 0.556 | 0.944 | 0.561 | 0.963 | 0.318 | 0.744 | 0.532 | 0.853 | ||
| GPT-4Par | 0.562 | 1.000 | 0.576 | 0.971 | 0.618 | 0.976 | 0.532 | 0.925 | 0.466 | 0.904 | 0.571 | 0.864 | ||
| Model | Tweet | GoogleNewsT | StackOverFlow | |||||||||||
| DC | Fa | DC | Fa | DC | Fa | DC | Fa | DC | Fa | DC | Fa | |||
| LDA | 0.337 | 1.000 | 0.561 | 1.000 | 0.488 | 1.000 | 0.664 | 1.000 | 0.684 | 1.000 | 0.842 | 1.000 | ||
| LDAAug | 0.307 | 1.000 | 0.579 | 0.997 | 0.531 | 1.000 | 0.763 | 1.000 | 0.659 | 1.000 | 0.838 | 1.000 | ||
| TSCTM | 0.176 | 1.000 | 0.388 | 1.000 | 0.405 | 1.000 | 0.740 | 1.000 | 0.141 | 1.000 | 0.480 | 1.000 | ||
| TSCTMAug | 0.187 | 1.000 | 0.331 | 0.987 | 0.309 | 1.000 | 0.608 | 0.979 | 0.419 | 0.888 | 0.441 | 0.888 | ||
| BERTopic | 0.293 | 1.000 | 0.471 | 1.000 | 0.433 | 1.000 | 0.748 | 1.000 | 0.656 | 1.000 | 0.796 | 1.000 | ||
| BERTopicAug | 0.303 | 1.000 | 0.468 | 1.000 | 0.422 | 1.000 | 0.749 | 1.000 | 0.637 | 1.000 | 0.795 | 1.000 | ||
| GPT-3.5Seq | 0.197 | 0.984 | 0.335 | 0.967 | 0.334 | 0.975 | 0.583 | 0.954 | 0.479 | 1.000 | 0.689 | 0.994 | ||
| GPT-3.5Par | 0.213 | 1.000 | 0.384 | 0.994 | 0.321 | 0.968 | 0.585 | 0.952 | 0.636 | 1.000 | 0.694 | 1.000 | ||
| GPT-4Seq | 0.224 | 0.983 | 0.403 | 0.994 | 0.373 | 1.000 | 0.660 | 0.951 | 0.554 | 0.931 | 0.626 | 0.883 | ||
| GPT-4Par | 0.241 | 1.000 | 0.402 | 1.000 | 0.392 | 1.000 | 0.661 | 0.995 | 0.578 | 1.000 | 0.754 | 1.000 | ||
4 Experiments
We conduct experiments to investigate how well our methods perform topic modeling for short texts.
4.1 Dataset
We employ three tokenized datasets provided by Zhang et al. (2021): GoogleNewsT Rakib et al. (2020), Tweet Yin and Wang (2016), and StackOverFlow.111https://www.kaggle.com/competitions/predict-closed-questions-on-stack-overflow/data?select=train.zip Following Wu et al. (2022), the datasets are preprocessed as follows: () characters are converted to lower case; () words of two letters or less are removed; () words appearing less than five times are filtered out. Then, we split each preprocessed dataset into subsets for our methods, setting the size at 1000222In preliminary experiments, we checked the performance of our methods under the setting where the subset size is 250, 500, and 1000. See Appendix A.3. and truncating the remaining example. Table 2 shows the final statistics of the datasets we use. Note that baseline models take the union of subsets as input, and each subset contains different examples for each run.
4.2 Model
In the study, we evaluate our approaches with GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-0125-preview) provided by the OpenAI API.333In preliminary experiments, we also tried Llama 2 Touvron et al. (2023), but found it was not controllable enough for its output to be applied to our approach. See Appendix A.2. As baseline models, we employ three models mentioned in Section 2: LDA444https://github.com/BobXWu/TopMost, TSCTM44footnotemark: 4, and BERTopic.555https://maartengr.github.io/BERTopic Additionally, we report the results of each baseline model with data augmentation. Regarding data augmentation techniques and hyperparameters of TSCTM, we followed the original settings Wu et al. (2022).666The details can be found in Appendix B.1.
4.3 Evaluation
We evaluate models under the conditions that the number of topics is 5 or 15, and the number of each topic words is 5. For evaluation metrics, we use two widely used metrics for topic quality and two new metrics to consider possible issues of LLMs, i.e., the possibility of outputting topics reflecting only very limited documents or ones not included in documents. We run each model five times and report the average scores.
Topic Coherence and Diversity
Following Wu et al. (2022), we calculate a coherence value777https://github.com/dice-group/Palmetto (Cv, Röder et al., 2015) with Wikipedia for topic coherence, and a topic uniqueness (TU, Nan et al., 2019) to evaluate the diversity in the inferred topics.
Document Coverage
We are concerned that LLMs infer topics that reflect only very limited documents. Thus we propose a metric document coverage, which measures the extent to which discovered topics cover documents. Document coverage is defined as follows:
where is the set of words present in each document . and denote the original document set without augmentation and the word set of the -th topic, respectively. A higher DC means that discovered topics cover more documents.
Factuality
Another potential issue is hallucination, where topics discovered by LLMs may not be included in given documents. Therefore we introduce factuality, which measures the degree to which topic words are composed from the vocabulary in original documents. Factuality is defined as follows:
A higher Fa indicates that more topic words are extracted from the vocabulary in the documents.
5 Results and Discussion
Topic Quality
Table 3 shows that the topics discovered by our methods are relatively high-quality both in terms of coherence (Cv) and diversity (TU).888We show examples of topics in Appendix C.1. In particular, for coherence, GPT-4 achieved the state-of-the-art performance in all settings, with up to 40 % improvement. For instance, on GoogleNewsT, the scores have risen by 41% (from 0.439 to 0.618) and 28% (from 0.437 to 0.561), respectively, for each setting of the number of topics.
Document Coverage
Table 4 reports that LLMs show relatively lower scores for document coverage (DC) than the best baseline models. This means that the topics discovered by LLMs often cover fewer documents than those discovered by the baseline models. However, note that there is a trade-off between topic coherence (Cv) and document coverage. For example, LDA achieved the highest coverage on GoogleNewsT but showed the lowest coherence, with the exception of TSCTM and TSCTM.
Factuality
In Table 4, LLMs show lower scores for factuality (Fa) than the baseline models, particularly ones without augmentation. This indicates that some topic words output by LLMs are not included in the documents. However, their factuality loss is less than 5% in almost all settings. Furthermore, we analyze these not-existent words and find that most are not too problematic to mislead topic interpretation; they are synonyms, derivatives, and related words of ones in the documents.999We show examples of not-existent words and analysis details in Appendix C.2. This suggests that LLMs do not generate hallucinated topics that cause misinterpretation of the content.
Parallel and Sequential Prompting
Table 3 and Table 4 show that the parallel prompting approach can identify topics with better coherence and document coverage than the sequential prompting one. To analyze the superior performance of the parallel approach, we calculate Cv and DC of topics before merging. Table 5 shows that Cv and DC scores before merging degrade from those of the parallel approach, demonstrating that the merging process can improve both their coherence and document coverage. On the other hand, we analyze the transition of topics during sequential approach and then observe that it tends to update the previously identified topic very little due to strict adherence to our instruction, leading to lower document coverage compared to the parallel approach.101010We show examples of their processing and further analysis in Appendix C.3.
| Model | Tweet | GoogleNewsT | StackOverFlow | |||||
| Cv | DC | Cv | DC | Cv | DC | |||
| GPT-3.5 | 0.532 | 0.366 | 0.517 | 0.569 | 0.464 | 0.634 | ||
| GPT-4 | 0.580 | 0.395 | 0.523 | 0.665 | 0.519 | 0.747 | ||
6 Conclusion
In this study, we propose two approaches to use LLMs for topic modeling: parallel prompting and sequential prompting. We evaluated the performance of GPT-3.5 and GPT-4 with our approaches compared to existing topic models on three datasets. In the evaluation, in addition to the well-known metrics for topic quality, we introduced two new metrics, document coverage and factuality, to assess the potential issues about LLMs reflecting only some documents or outputting hallucinated topics. The results showed that LLMs could find higher-quality topics than existing methods, and the impact of the potential issues was not remarkable in practice. Future work includes improving our methods to make it possible to assign a topic to each document.
Acknowledgements
This work was supported by JSPS KAKENHI Grant Number JP24H00809, Japan.
Appendix A Preliminary Experiments
In preliminary experiments, we search prompts and a subset size for our methods to maximize the performance of our methods.
A.1 Prompts
We first considered the Par prompt and then proceeded to the Par prompt and the Seq prompt.
Par
We checked three kinds of prompts, as shown in Table 6. We finally selected a Direct prompt as a tentative a Par prompt, which achieved the highest performance We then also consider the effect of the insertion of the following phrases, which are expected to improve scores of topic coherence, diversity, and document coverage, respectively.
- Cv
“NOTE: Make top words for each topic likely to occur together in the documents”
- TU
“NOTE: Make the top words unique across topics.”
- DC
“NOTE: Maximize the number of documents that contain at least one of the top words.”
However, we found that none of the above can positively influence LLMs’ performance in our methods. Therefore, we selected a Direct prompt without phrase insertion as the Par prompt.
Par
Considering the Par prompt, we created a Base Par prompt, which has a similar structure to the Par (Table 7). Then, we consider the insertion of the following phrases:
- Goal
“We aim to identify topics for the entire document set by merging the topic modeling results for each subset.”
- Detail
“NOTE: Outputs should reflect the topics before merging as much as possible. Output should contain topics that often appear before merging and not have ones that don’t appear much before merging.”
Experimental results showed our methods performed the best when we inserted both the Goal phrase and the Detail phrase into the Base Par.
As a result, we employ a Base Par prompt with both phrases as the Par prompt for the parallel approach.
Seq
Similar to the prompt for parallel, we first created a simple Base Seq prompt for the sequential approach in Table 7, and validated the effect of the insertion of the following phrases.
- Goal
“We aim to identify topics for the entire document set by sequentially updating tentative topics identified from each subset, considering topics identified just before from another subset.”
- Detail
“NOTE: Outputs should be the same as the previous topics as much as possible. You can change them minimally only when the given documents don’t include them much, and a new topic needs to be added to describe the documents.”
We also found that the insertion of both above phrases is most effective in improving the performance of the sequential method. Thus, we utilize a Base Seq prompt, incorporating both phrases as the Seq prompt for the sequential approach.
A.2 Llama 2
In preliminary experiments, we also tried llama-2-7b-chat and llama-2-13b-chat as LLMs for our methods and found that it is difficult for Llama 2 Touvron et al. (2023) to perform topic modeling regardless of prompts and the subset size we use. Table 8 shows outputs of Llama 2 when given the Par prompt with the subset size 100 on GoogleNewsT. Llama 2 can not make adequate output following our instruction for the number of topics and topic words, while GPT-3.5 and GPT-4 can perform it steadily under the same setting.
A.3 Subset Size
We assume candidates of the subset size are 250, 500, and 1000. It would be difficult to adopt more than 1000 as the subset size due to the context length of GPT-3.5 (gpt-3.5-turbo-0125), which we plan to use for the main experiments.
We run the parallel and the sequential methods with GPT-3.5 on GoogleNewsT in each case of the subset size. Table 9 reports the average scores of each method for five runs. We found the tendency of improving topic coherence as the subset size increase, but could not find any tendency for the other metrics. We finally select 1000 as the subset size because the performance of each models are relatively high in all metrics under that setting.
Appendix B Experimental Details
B.1 Implementation Details
We run TSCTM for 200 epochs. In the case without data augmentation, we run it with temperature as 0.5 and weight contrast as 1.0. In the case of data augmentation, we run it with temperature as 0.07, weight contrast as 3.0, and same quant as 0.001. For data augmentation, we applied WordNet111111https://github.com/makcedward/nlpaug and Contextual Augmenter33footnotemark: 3 Kobayashi (2018) with 30% word replacement, and filtered low-frequency words as in the preprocessing. Each Augmenter randomly replaces words in an input text with synonyms defined by WordNet and with words predicted by BERT Devlin et al. (2019)121212https://huggingface.co/bert-base-uncased, respectively. For gpt-3.5-turbo-0125, gpt-4-0125-preview, and BERTopic, we utilized their original configurations without modification.
B.2 Examples of Prompts
Table 10 shows examples of prompts used in the experiment.
Appendix C Result Details
C.1 Examples of Topics
Following Wu et al. (2022), we randomly select some examples of topics identified by LDA, BERTopic, and our proposed methods with GPT-4.
C.2 Examples of Topic Words Not Included in the Documents
Table 12 shows examples of words not included in the documents outputted in topic modeling on GoogleNewsT. GPT-3.5 output in bold are the names of entities (e.g., broncos, gree, and watson) or words that do not exist in the real world (e.g., dorffiefskee).
Such words are considered harmful because they may induce misinterpretation of topics. However, only a small number of such words were found, and most of them were synonyms, synonyms, derivatives, or related words in the documents.
C.3 Examples of the Processing
Table C.3 reports the concrete examples of topics identified for each subset and the final output to demonstrate the processing in our methods. In the parallel approach, we find that LLM reasonably merges topics from each subset. For instance, bold topics in each subset are merged into one topic in the final output, using words from both subsets. On the other hand, in the sequential approach, the final output is the same as the topics for the first subset except for only one pair of bold words. This indicates that LLMs with the the sequential approach could too strictly keep topics for the previous subset so that they can not output topics that sufficiently reflect the entire set.
Appendix D Limitations
We do not consider it matters whether GPT-3.5 and GPT-4 have seen these datasets in the pre-training; since topic modeling is an unsupervised task, they can not utilize their knowledge about these datasets in our experiment. Our approaches perform promisingly in a practical common situation where the number of topics is about 10 but are less reliable with a larger number of topics, such as 50 or 100, due to the difficulty of controlling the LLM’s output format.
| ID | Candidates of Base Prompt Template |
| Direct | Write the results of simulating topic modeling for the following documents, each starting with "#." Assume you will finally identify [NUM_TOPICS] topics and use 5 top words for each topic. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. """ [DOCS] """ |
| Indirect | Discover latent [NUM_TOPICS] topics in the following documents, each starting with "#." For each topic, write 5 words extracted from input texts to show its meanings. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. """ [DOCS] """ |
| Direct | """ [DOCS] """ Write the results of simulating topic modeling for the above documents, each starting with "#." Assume you will finally identify [NUM_TOPICS] topics and use 5 top words for each topic. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. |
| ID | Candidates of Base Prompt Template |
| Base Par | Write the results of merging the following topic modeling results for each subset of the document set. Each result starts with "- n" and its topics start with "#" """ - 1 [TOPICS] - 2 [TOPICS] - 3 … """ |
| Base Seq | Write the results of simulating topic modeling for the following documents, each starting with "#." Make the most use of the following topics previously identified from another set of documents, each starting with "Topic k:": """ [TOPICS] """ Assume you will finally identify [NUM_TOPICS] topics and use 5 top words for each topic. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. """ [DOCS] """ |
| Model size | Examples of Llama2 Output |
| 7B | Topic 1: Top words: relief, challenge, face Topic 2: Top words: welker, concussion, test Topic 3: Top words: live, stream, champion, league Topic 4: Top words: bargain, black, friday, shopping Topic 5: Top words: scotland, independence, white, paper Note: Each topic is represented by 5 top words, which are the most frequently occurring words in the given documents. |
| 13B | Topic 1: Disasters and Relief Efforts Topic 2: Sports and Injuries Topic 3: Technology and Gadgets Topic 4: Politics and Leadership Topic 5: Entertainment and Celebrities |
| Subset Size | Cv | TU | DC | Fa |
| 0.531 | 0.936 | 0.241 | 1.000 | |
| 0.572 | 0.896 | 0.241 | 1.000 | |
| 0.571 | 0.960 | 0.213 | 1.000 |
| Subset Size | Cv | TU | DC | Fa |
| 0.524 | 0.976 | 0.198 | 0.992 | |
| 0.529 | 0.992 | 0.193 | 0.976 | |
| 0.562 | 0.984 | 0.197 | 0.984 |
| ID | Prompt Example |
| Par | Write the results of simulating topic modeling for the following documents, each starting with "#." Assume you will identify 5 topics and use 5 top words for each topic. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. """ # philippine typhoon relief effort face challenge # wes welker concussion test bronco # basel chelsea live stream champion league watch … # discus black friday shopping secret """ |
| Par |
We aim to identify topics for the entire document set by merging the topic modeling results for each subset.
Write the results of merging the following topic modeling results for each subset of the document set.
Each result starts with "- n" and its topics start with "#" """ - 1 # comet ison thanksgiving sun solar # kanye west bound parody video # nokia lumia release mobile device # black friday shopping thanksgiving sale # alec baldwin msnbc cancellation defends … - 11 # nokia lumia sale december phone # kanye west kim kardashian taylor # black friday deal best sales # irs rule political activity tax # bronco patriot win game rivalry """ Assume you will finally identify 5 topics and use 5 top words for each topic. NOTE: Outputs should reflect the topics before merging as much as possible. Output should contain topics that often appear before merging and not have ones that don’t appear much before merging. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. |
| Seq |
We aim to identify topics for the entire document set by sequentially updating tentative topics identified from each subset, considering topics identified just before from another subset.
Write the results of simulating topic modeling for the following documents, each starting with "#."
Make the most use of the following topics previously identified from another set of documents, each starting with "Topic k:":
""" Topic 1: kanye west kim kardashian bound Topic 2: xbox black friday cyber monday Topic 3: hewlett packard nokia lumia company Topic 4: dancing star finale winner season Topic 5: syria peace talk china air """ Assume you will finally identify 5 topics and use 5 top words for each topic. NOTE: Outputs should be the same as the previous topics as much as possible. You can change them minimally only when the given documents don’t include them much, and a new topic needs to be added to describe the documents. NOTE: Outputs must always be in the format "Topic k: word word word word word" and nothing else. """ # spacex falcon launch attempt # taylor swift princess gown winter white # redbox instant window phone appears nokia exclusive … # google backed company selling dna analysis kit ordered sale """ |
| Model | Examples of topics |
| LDA | xbox microsoft game patriot bronco |
| nokia lumia oldboy launch google | |
| kobe bryant chelsea lakers basel | |
| TSCTM | macy parade hanukkah thanksgiving travel |
| china zone african japan johansson | |
| bronco patriot packer welker illinois | |
| BERTopic | china zone air nsa porn |
| methane ant emission fire burning | |
| thanksgiving friday black comet parade | |
| GPT-4 | wes welker nfl concussion game |
| nokia lumia window phone december | |
| nfl season game player concussion | |
| GPT-4 | san andreas mobile game release |
| nokia lumia tablet smartphone launch | |
| thanksgivukkah hanukkah holiday feast rare |
| Model | Examples of the topic words not included in the documents |
| TACTM | twelvemonth sink railway blowout |
| GPT-3.5 | dorffiefskee broncos patriots health advancement ocean guilty france legal attorney |
| GPT-3.5 | gree watson advertisement boat funding attorney declared refugees crash digital |
| Subset 1 |
| fishing fish bass fly report |
| superbowl commercial bowl super best |
| king speech oscar nomination award |
| facebook privacy setting user change |
| acai berry weight loss diet plan |
| Subset 2 |
| fishing fish fly book saltwater |
| superbowl commercial doritos pepsi volkswagen |
| king speech oscar nomination award best |
| acai berry weight loss diet plan |
| christina aguilera national anthem super bowl |
| Final Output |
| fishing fish fly bass saltwater |
| superbowl commercial bowl pepsi doritos |
| king speech oscar nomination award |
| acai berry weight loss diet health |
| facebook privacy setting user change |
| Subset 1 |
| fishing commercial superbowl fly bass |
| facebook privacy setting user setting |
| king speech oscar nomination award |
| berry acai weight diet loss |
| christina aguilera national anthem super |
| Final Output |
| fishing fly superbowl commercial bass |
| facebook privacy setting user security |
| king speech oscar nomination award |
| acai berry weight diet loss |
| christina aguilera national anthem super |

### v2

@@
-Topic Modeling for Short Texts with Large Language Models
+Comprehensive Evaluation of Large Language Models for Topic Modeling
 Abstract
-As conventional topic models rely on word co-occurrence to infer latent topics, topic modeling for short texts has been a long-standing challenge.
-Large Language Models (LLMs) can potentially overcome this challenge by contextually learning the semantics of words via pretraining.
-This paper studies two approaches, parallel prompting and sequential prompting, to use LLMs for topic modeling.
-Due to the input length limitations, LLMs cannot process many texts at once.
-By splitting the texts into smaller subsets and processing them parallelly or sequentially, an arbitrary number of texts can be handled by LLMs.
-Experimental results demonstrated that our methods can identify more coherent topics than existing ones while maintaining the diversity of the induced topics.
-Furthermore, we found that the inferred topics adequately covered the input texts, while hallucinated topics were hardly generated.
-Topic Modeling for Short Texts with Large Language Models
+Recent work utilizes Large Language Models (LLMs) for topic modeling, generating comprehensible topic labels for given documents.
+However, their performance has mainly been evaluated qualitatively, and there remains room for quantitative investigation of their capabilities.
+In this paper, we quantitatively evaluate LLMs from multiple perspectives: the quality of topics, the impact of LLM-specific concerns, such as hallucination and shortcuts for limited documents, and LLMs’ controllability of topic categories via prompts.
+Our findings show that LLMs can identify coherent and diverse topics with few hallucinations but may take shortcuts by focusing only on parts of documents.
+We also found that their controllability is limited.
+Comprehensive Evaluation of Large Language Models for Topic Modeling
 Tomoki Doi The University of Tokyo doi-tomoki701@is.s.u-tokyo.ac.jp Masaru Isonuma The University of Edinburgh m.isonuma@ed.ac.uk Hitomi Yanaka The University of Tokyo hyanaka@is.s.u-tokyo.ac.jp
 1 Introduction
+Recent studies have introduced topic modeling using prompt-based LLMs Ouyang et al.
+(2022); OpenAI (2023), which directly generate comprehensible topic descriptions in given documents.
+Specifically, Pham et al.
+(2023) shows that topics generated by LLMs are highly consistent with human-annotated topic labels, and Mu et al.
+(2024) demonstrates that well-crafted prompts could control the granularity of the generated topics.
+However, since previous studies primarily evaluate the quality of topics through human judgment, it remains unclear to what extent the topics generated by LLMs meet the criterion of coherence and diversity considered in the conventional topic modeling.
+In addition, LLM-specific concerns have not been considered enough in the previous evaluations; it is not impractical to manually examine the impact of hallucinations and shortcuts, in which LLMs reflect only parts of documents.
+LLMs may generate topics that induce misinterpretation of documents due to these issues, and they may not necessarily be superior to conventional topic models.
+In this paper, we quantitatively evaluate LLMs on topic modeling from multiple perspectives: the quality of generated topics, the risk of hallucinations and shortcuts, and their controllability.
+We conduct experiments employing two prompting methods on three datasets from various domains.
+To assess the quality of topics generated by LLMs, we directly compare LLMs with conventional topic models using well-established metrics: topic coherence and diversity.
+Additionally, we introduce document coverage and factuality to assess the impact of hallucinations and shortcuts; document coverage indicates how many documents are covered by the generated topics, and factuality evaluates the extent to which the topics exist in the documents.
+Furthermore, we explore the controllability of LLMs by examining how effectively the category of the generated topics can be controlled through prompts.
+The contributions of this study are as follows:
+1.
+We validate the performance of LLMs for topic modeling by comparing them with conventional models and show that LLMs can identify more coherent topics while maintaining the diversity of topics.
+2.
+We assess the document coverage and factuality of topics generated from LLMs.
+Experiments indicate few issues with hallucination but suggest the possibility of taking shortcuts by focusing only on parts of given documents.
+3.
+We examine the controllability of LLMs in generating topics associated with a predefined category specified in prompts, quantitatively demonstrating their limitations.
+2 Related Work
 Topic modeling is a classical task of discovering latent topics that best describe a set of documents Blei et al.
 (2003); Churchill and Singh (2022).
-Recently, while neural topic models work successfully on various kinds of long documents Miao et al.
-(2017); Srivastava and Sutton (2017); Dieng et al.
-(2020a), they can not handle short texts, such as social media posts and news headlines Li et al.
-(2016); Wu et al.
+While traditional latent Dirichlet allocation (LDA, Blei et al., 2003) was the primary approach, neural topic models and clustering-based methods using contextual embeddings have shown impressive performance Srivastava and Sutton (2017); Dieng et al.
+(2020); Zhang et al.
 (2022).
-Large Language Models (LLMs), such as InstructGPT Ouyang et al.
-(2022) and GPT-4 OpenAI (2023), have shown impressive results on various tasks by providing task instructions in a zero-shot manner (Wang et al., 2023; Kocoń et al., 2023).
-Since conventional topic models infer the topics of words by relying on word co-occurrence, they perform worse on short texts.
-In contrast, as LLMs contextually learn the semantics of words by pretraining on massive text corpora, they could accurately infer the latent topics of short texts.
-In order to use LLMs for topic modeling, we propose two approaches, parallel prompting and sequential prompting (Figure 1).
-Due to the input length limitations of LLMs, an input document set must be split into smaller subsets, which are processed individually.
-Parallel prompting concurrently infers the topics of each subset and merges them to represent the topics of the whole document set.
-Sequential prompting processes each subset successively, updating the topics in every iteration.
-We assess our approaches across texts from various domains using multiple evaluation metrics.
-The contributions of this study are as follows:
-1.
-We propose parallel and sequential prompting methods for topic modeling using LLMs.
-Our methods can handle a large number of texts that cannot be processed in a single run due to the input length limitations of LLMs.
-2.
-We validate the performance of our methods by comparing them to existing models and show that ours can identify more coherent topics than existing models while maintaining the diversity of the induced topics.
-3.
-We assess the document coverage and factuality of the inferred topics, due to concerns that LLMs may focus on only parts of documents or generate hallucinated topics.
-Evaluation results indicate that our concerns are negligible.
-2 Background
-Topic modeling is a task of identifying latent topics as a set of topic words representing each topic from a collection of documents Blei et al.
+Recent work Pham et al.
+(2023); Mu et al.
+(2024) has introduced prompt-based approaches using LLMs for topic modeling, which generates labels or brief sentences as topics, instead of conventional topic words (i.e., a set of words in documents).
+To leverage input length constraints of LLMs, Mu et al.
+(2024) uses a parallel strategy, where LLMs identify topics for each document and then merge them.
+Pham et al.
+(2023) employs a sequential strategy that supplies LLMs with each document along with previously identified topics and iteratively updates the topic list.
+These studies primarily evaluated topics generated by LLMs through human judgment based on topic visualization Mu et al.
+(2024) and ground-truth labels Pham et al.
 (2023).
-Topic modeling has conventionally been tackled with probabilistic models such as latent Dirichlet allocation (LDA, Blei et al., 2003).
-In recent years, however, neural models have become popular due to their high performance Srivastava and Sutton (2017); Dieng et al.
-(2020b); Grootendorst (2022).
-It is known that topic modeling for short texts is difficult for current topic models due to data sparsity Li et al.
-(2016); Wu et al.
-(2022).
-TSCTM Wu et al.
-(2022) is a current state-of-the-art neural topic model for short texts.
-This model addresses data sparsity by learning representations of documents based on VQ-VAE van den Oord et al.
-(2017), contrastive learning, and incorporation of data augmentation into the learning.
-BERTopic Grootendorst (2022) uses a pretrained encoder, Sentence-BERT Reimers and Gurevych (2019), to obtain clusters of documents and assigns topic words to each cluster by a class-based TF-IDF procedure.
-Another related study is Stammbach et al.
-(2023), which utilizes LLMs to evaluate topic quality automatically.
-However, our study is the first to explore how well LLMs perform topic modeling.
+It has been indicated that they could better align with human preferences compared to conventional topic models.
+However, previous studies do not consider quantitative assessment of the coherence and diversity of topics.
+In addition, they do not investigate LLM issues, such as hallucination and shortcuts.
 3 Topic Modeling with LLMs
-We introduce two approaches to perform topic modeling with LLMs: parallel prompting and sequential prompting.
-For these approaches, we apply common preprocessing, which involves randomly splitting a document set into subsets with the same size, smaller than the context length of the LLMs.
+Inspired by the previous work Pham et al.
+(2023); Mu et al.
+(2024), we consider two approaches to perform topic modeling with LLMs: parallel prompting and sequential prompting.
+We use the three prompts shown in Table 1.
+For these approaches, we apply common preprocessing, which involves randomly splitting a document set into subsets of the same size, smaller than the context length of the LLMs.
+To facilitate detailed evaluations, we adopt topic word representation for generated topics.
 Parallel Prompting
-In the parallel prompting, LLMs identify topics for each subset in parallel by prompting the subset and the instruction of topic modeling.
-Then, the topics of each subset are merged by LLMs.
-We use two kinds of prompts shown in Table 1: (i) a Par prompt for parallel topic modeling for each subset, and (ii) a Par prompt for merging topics of each results.
+In the parallel prompting, LLMs with Par prompt execute topic modeling for each subset in parallel. Then, the topics of each subset are merged by LLMs with Par prompt.
 Sequential Prompting
-In the sequential prompting, LLMs identify topics for each subset sequentially, considering the topics previously identified for the previous subset.
-We use a Par for the first subset, then use a Seq prompt in Table 1 for the other subsets.
-This prompt contains topics identified in the prior subset and instructions to refer to them.
+In the sequential prompting, LLMs identify topics for each subset sequentially, considering the previously identified topics.
+We use Par prompt for the first subset, then use Seq prompt for the other subsets.
 | ID | Prompt |
@@
 | for the following documents: [DOCS], Make the | |
-| most use of the following topics: [TOPICS]. |
-| # of | Text | Vocabulary | |
-| Dataset | Documents | Length | Size |
-| Tweet | 2000 | 5.47 | 706 |
-| GoogleNewsT | 11000 | 5.25 | 2376 |
-| StackOverFlow | 19000 | 4.71 | 2544 |
+| most use of the following topics: [TOPICS]. | |
+| Base/Orcl | Discover latent [K] topics (specifically related to |
+| (Ctrl) | [CAT]) in the following documents: [DOCS]. |
 | Model | Tweet | GoogleNewsT | StackOverFlow | |||||||||||
@@
 | BERTopicAug | 0.535 | 1.000 | 0.526 | 1.000 | 0.412 | 1.000 | 0.417 | 1.000 | 0.460 | 1.000 | 0.489 | 0.955 | ||
+| GPT-3.5Par | 0.476 | 0.992 | 0.532 | 0.900 | 0.571 | 0.960 | 0.535 | 0.913 | 0.312 | 0.864 | 0.496 | 0.913 | ||
 | GPT-3.5Seq | 0.552 | 0.960 | 0.515 | 0.920 | 0.562 | 0.984 | 0.489 | 0.948 | 0.441 | 0.896 | 0.517 | 0.775 | ||
-| GPT-3.5Par | 0.476 | 0.992 | 0.532 | 0.900 | 0.571 | 0.960 | 0.535 | 0.913 | 0.312 | 0.864 | 0.496 | 0.913 | ||
+| GPT-4Par | 0.562 | 1.000 | 0.576 | 0.971 | 0.618 | 0.976 | 0.532 | 0.925 | 0.466 | 0.904 | 0.571 | 0.864 | ||
 | GPT-4Seq | 0.577 | 0.992 | 0.551 | 0.976 | 0.556 | 0.944 | 0.561 | 0.963 | 0.318 | 0.744 | 0.532 | 0.853 | ||
-| GPT-4Par | 0.562 | 1.000 | 0.576 | 0.971 | 0.618 | 0.976 | 0.532 | 0.925 | 0.466 | 0.904 | 0.571 | 0.864 | ||
 | Model | Tweet | GoogleNewsT | StackOverFlow | |||||||||||
@@
 | BERTopicAug | 0.303 | 1.000 | 0.468 | 1.000 | 0.422 | 1.000 | 0.749 | 1.000 | 0.637 | 1.000 | 0.795 | 1.000 | ||
+| GPT-3.5Par | 0.213 | 1.000 | 0.384 | 0.994 | 0.321 | 0.968 | 0.585 | 0.952 | 0.636 | 1.000 | 0.694 | 1.000 | ||
 | GPT-3.5Seq | 0.197 | 0.984 | 0.335 | 0.967 | 0.334 | 0.975 | 0.583 | 0.954 | 0.479 | 1.000 | 0.689 | 0.994 | ||
-| GPT-3.5Par | 0.213 | 1.000 | 0.384 | 0.994 | 0.321 | 0.968 | 0.585 | 0.952 | 0.636 | 1.000 | 0.694 | 1.000 | ||
+| GPT-4Par | 0.241 | 1.000 | 0.402 | 1.000 | 0.392 | 1.000 | 0.661 | 0.995 | 0.578 | 1.000 | 0.754 | 1.000 | ||
 | GPT-4Seq | 0.224 | 0.983 | 0.403 | 0.994 | 0.373 | 1.000 | 0.660 | 0.951 | 0.554 | 0.931 | 0.626 | 0.883 | ||
-| GPT-4Par | 0.241 | 1.000 | 0.402 | 1.000 | 0.392 | 1.000 | 0.661 | 0.995 | 0.578 | 1.000 | 0.754 | 1.000 | ||
-4 Experiments
-We conduct experiments to investigate how well our methods perform topic modeling for short texts.
+4 Comparison with Conventional Models
+We quantitatively compare LLMs to conventional topic models, assessing topic quality and LLM-specific issues like shortcuts and hallucinations.
 4.1 Dataset
@@
 (2021): GoogleNewsT Rakib et al.
-(2020), Tweet Yin and Wang (2016), and StackOverFlow.111https://www.kaggle.com/competitions/predict-closed-questions-on-stack-overflow/data?select=train.zip Following Wu et al.
-(2022), the datasets are preprocessed as follows: () characters are converted to lower case; () words of two letters or less are removed; () words appearing less than five times are filtered out.
-Then, we split each preprocessed dataset into subsets for our methods, setting the size at 1000222In preliminary experiments, we checked the performance of our methods under the setting where the subset size is 250, 500, and 1000.
+(2020), Tweet Yin and Wang (2016), and StackOverFlow.111https://www.kaggle.com/competitions/predict-closed-questions-on-stack-overflow/data?select=train.zip First, the datasets are preprocessed using the preprocessing pipelines by Wu et al.
+(2022).
+Then, we split each preprocessed dataset into subsets for parallel and sequential prompting, setting the size at 1000222In preliminary experiments, we checked the performance of LLMs under the setting where the subset size is 250, 500, and 1000.
 See Appendix A.2.
-and truncating the remaining example.
-Table 2 shows the final statistics of the datasets we use.
-Note that baseline models take the union of subsets as input, and each subset contains different examples for each run.
+and truncating the remaining examples.333The preprocessing details and the final statistics of the datasets used in this study can be found in Appendix B.1 Note that models other than LLMs take the union of subsets as input.
 4.2 Model
-In the study, we evaluate our approaches with GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-0125-preview) provided by the OpenAI API.333In preliminary experiments, we also tried Llama 2 Touvron et al.
-(2023), but found it was not controllable enough for its output to be applied to our approach.
+In this experiment, we evaluate GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-0125-preview) provided by the OpenAI API as representative LLMs.444In preliminary experiments, we also tried Llama 2 Touvron et al.
+(2023), but found it was not controllable enough for its output to be applied to topic modeling.
 See Appendix A.3.
-As baseline models, we employ three models mentioned in Section 2: LDA444https://github.com/BobXWu/TopMost, TSCTM44footnotemark: 4, and BERTopic.555https://maartengr.github.io/BERTopic Additionally, we report the results of each baseline model with data augmentation.
-Regarding data augmentation techniques and hyperparameters of TSCTM, we followed the original settings Wu et al.
-(2022).666The details can be found in Appendix B.1.
+For baseline models, we consider three conventional topic models from different paradigms: LDA,555https://github.com/BobXWu/TopMost a state-of-the-art neural topic model (TSCTM,55footnotemark: 5 Wu et al.
+(2022)), and the popular clustering-based model (BERTopic,666https://maartengr.github.io/BERTopic Grootendorst (2022)).
+Additionally, we report the results of each model with data augmentation.
+All hyperparameters follow default implementations.
+777The details can be found in Appendix B.2.
 4.3 Evaluation
 We evaluate models under the conditions that the number of topics is 5 or 15, and the number of each topic words is 5.
-For evaluation metrics, we use two widely used metrics for topic quality and two new metrics to consider possible issues of LLMs, i.e., the possibility of outputting topics reflecting only very limited documents or ones not included in documents.
-We run each model five times and report the average scores.
+For evaluation metrics, we use two widely used metrics for topic quality and two new metrics to consider the issues of hallucination and shortcuts, i.e., the possibility of outputting topics not included in documents or reflecting only very limited documents.
+We run each model five times, each time with a different order of samples, and report the average score.
 Topic Coherence and Diversity
 Following Wu et al.
-(2022), we calculate a coherence value777https://github.com/dice-group/Palmetto (Cv, Röder et al., 2015) with Wikipedia for topic coherence, and a topic uniqueness (TU, Nan et al., 2019) to evaluate the diversity in the inferred topics.
+(2022), we calculate Cv888https://github.com/dice-group/Palmetto Röder et al.
+(2015) referencing Wikipedia for topic coherence, and TU Nan et al.
+(2019) to evaluate the diversity in the inferred topics.
 Document Coverage
-We are concerned that LLMs infer topics that reflect only very limited documents.
+We are concerned that LLMs take shortcuts, inferring topics that reflect only very limited documents.
 Thus we propose a metric document coverage, which measures the extent to which discovered topics cover documents.
 Document coverage is defined as follows:
-where is the set of words present in each document .
-and denote the original document set without augmentation and the word set of the -th topic, respectively.
-A higher DC means that discovered topics cover more documents.
+where is a document within the reference document collection, and is the topic word constituting the outputted topics.
+A higher DC means that discovered topics cover more reference documents.
+In this experiment, we use the preprocessed datasets without augmentation as references.
 Factuality
-Another potential issue is hallucination, where topics discovered by LLMs may not be included in given documents.
-Therefore we introduce factuality, which measures the degree to which topic words are composed from the vocabulary in original documents.
-Factuality is defined as follows:
-A higher Fa indicates that more topic words are extracted from the vocabulary in the documents.
-5 Results and Discussion
+Another potential issue is hallucination, where topics discovered by LLMs may not be included in given documents. Therefore, we introduce factuality as follows:
+A higher Fa indicates that more topic words are composed from the vocabulary in the reference documents.
+4.4 Results and Discussion
 Topic Quality
-Table 3 shows that the topics discovered by our methods are relatively high-quality both in terms of coherence (Cv) and diversity (TU).888We show examples of topics in Appendix C.1.
+Table 2 shows that the topics generated by LLMs are relatively high-quality both in terms of coherence (Cv) and diversity (TU).999We show examples of topics in Appendix C.1.
 In particular, for coherence, GPT-4 achieved the state-of-the-art performance in all settings, with up to 40 % improvement.
@@
 Table 3 reports that LLMs show relatively lower scores for document coverage (DC) than the best baseline models.
-This means that the topics discovered by LLMs often cover fewer documents than those discovered by the baseline models.
-However, note that there is a trade-off between topic coherence (Cv) and document coverage.
-For example, LDA achieved the highest coverage on GoogleNewsT but showed the lowest coherence, with the exception of TSCTM and TSCTM.
+This means that the topics generated by LLMs often cover fewer documents than those discovered by the conventional models, which suggests the presence of the shortcuts issue.
+Note that in both GPT-3.5 and GPT-4, parallel prompting results in higher DC scores compared to sequential prompting, implying it might help mitigate the shortcuts issue.101010Further analysis is provided in Appendix C.2.
 Factuality
 In Table 3, LLMs show lower scores for factuality (Fa) than the baseline models, particularly ones without augmentation.
-This indicates that some topic words output by LLMs are not included in the documents.
+This indicates that some topic words generated by LLMs are not included in the documents.
 However, their factuality loss is less than 5% in almost all settings.
-Furthermore, we analyze these not-existent words and find that most are not too problematic to mislead topic interpretation; they are synonyms, derivatives, and related words of ones in the documents.999We show examples of not-existent words and analysis details in Appendix C.2.
-This suggests that LLMs do not generate hallucinated topics that cause misinterpretation of the content.
-Parallel and Sequential Prompting
-Table 3 and Table 4 show that the parallel prompting approach can identify topics with better coherence and document coverage than the sequential prompting one.
-To analyze the superior performance of the parallel approach, we calculate Cv and DC of topics before merging.
-Table 5 shows that Cv and DC scores before merging degrade from those of the parallel approach, demonstrating that the merging process can improve both their coherence and document coverage.
-On the other hand, we analyze the transition of topics during sequential approach and then observe that it tends to update the previously identified topic very little due to strict adherence to our instruction, leading to lower document coverage compared to the parallel approach.101010We show examples of their processing and further analysis in Appendix C.3.
-| Model | Tweet | GoogleNewsT | StackOverFlow | |||||
-| Cv | DC | Cv | DC | Cv | DC | |||
-| GPT-3.5 | 0.532 | 0.366 | 0.517 | 0.569 | 0.464 | 0.634 | ||
-| GPT-4 | 0.580 | 0.395 | 0.523 | 0.665 | 0.519 | 0.747 | ||
+Furthermore, we analyze these non-existent words and find that they are mostly synonyms, derivatives, and related words of words in the documents.111111We show examples of not-existent words and analysis details in Appendix C.3.
+This suggests that LLMs might not frequently generate hallucinated topics that could lead to misinterpretation of the content.
+5 Controlled Topic Modeling
+In this section, we investigate whether specifying the category of topics via prompts can effectively control the category of the generated topics.
+121212Note that we also examine the controllability of the number of topics, demonstrating that LLMs can flexibly generate as many topics as specified in prompts.
+See Appendix D.
+Settings
+We employ the widely-used 20 NewsGroups131313http://qwone.com/~jason/20Newsgroups/ dataset annotated with six major topic categories.
+We randomly extract 500 samples and apply the same preprocessing as in Section 4.1.
+We evaluate topic modeling with GPT-4 (gpt-4-turbo-2024-04-09) under three different prompt settings: Ctrl, Base, and Orcl, as shown in Table 1.
+Using Ctrl prompt, GPT-4 conducts controlled topic modeling focused on the category specified in [CAT]: computer, politics, or science.
+Under Base prompt, GPT-4 executes regular topic modeling without any category specification.
+Furthermore, as a hypothetical upper bound on the performance with Ctrl prompt, we introduce an oracle model with Orcl prompt, which performs topic modeling provided with only documents annotated with [CAT].
+Results and Discussion
+To assess the controllability of GPT-4 with Ctrl prompt, we calculate document coverage (DC; Section 4.3), using documents in the [CAT] category as reference documents.
+Table 4 demonstrates that GPT-4 achieves higher DC scores with Ctrl prompt compared to Base prompt, but scores lower than GPT-4 with Orcl prompt.
+This suggests that LLMs may have limited controllability in identifying topics related to a specified category in prompts, indicating opportunities for improvement.
+Furthermore, to explore the influence of controlled topic modeling via prompts on hallucination issues, we calculate factuality (Fa; Section 4.3) using all provided samples as the reference.
+Table 4 shows that Fa of GPT-4 when using Ctrl prompt is lower than that when using Base prompt, implying that specifying a category in prompts may increase the hallucination risk.
+| Prompt | |||
+| Base | 0.621 | 0.530 | 0.651 |
+| Orcl | 0.968 | 0.823 | 0.884 |
+| Ctrl | 0.925 | 0.719 | 0.767 |
+| Prompt | |
+| Base | 0.968 |
+| Ctrl | 0.959 |
+| Ctrl | 0.936 |
+| Ctrl | 0.895 |
 6 Conclusion
-In this study, we propose two approaches to use LLMs for topic modeling: parallel prompting and sequential prompting.
-We evaluated the performance of GPT-3.5 and GPT-4 with our approaches compared to existing topic models on three datasets.
-In the evaluation, in addition to the well-known metrics for topic quality, we introduced two new metrics, document coverage and factuality, to assess the potential issues about LLMs reflecting only some documents or outputting hallucinated topics.
-The results showed that LLMs could find higher-quality topics than existing methods, and the impact of the potential issues was not remarkable in practice.
-Future work includes improving our methods to make it possible to assign a topic to each document.
-Acknowledgements
-This work was supported by JSPS KAKENHI Grant Number JP24H00809, Japan.
+This paper comprehensively evaluated LLMs for topic modeling, focusing on topic quality, concerns with LLMs, and their controllability in guiding generated topics via prompts.
+Our results show that LLMs can identify coherent topics without compromising diversity or facing significant hallucination issues, while they might take shortcuts by focusing only on parts of given documents.
+Furthermore, our study provides quantitative evidence indicating that the controllability of LLMs through prompts is limited.
+7 Limitations
+Since we evaluated LLMs using GPT-3.5 and GPT-4, both of which are closed-source models, we could not thoroughly consider whether their pre-training and instruction-tuning datasets might contain the datasets used in this study.
+Current open-source LLMs, such as Llama 2, do not sufficiently adhere to the output format needed for evaluation, as shown in Appendix A.3.
+We hope future work will evaluate more advanced open-source LLMs through our frameworks and offer more transparent evaluations.
 Appendix A Preliminary Experiments
-In preliminary experiments, we search prompts and a subset size for our methods to maximize the performance of our methods.
+In preliminary experiments, we search prompts and a subset size for our prompting methods to maximize the performance of LLMs.
 A.1 Prompts
@@
 Thus, we utilize a Base Seq prompt, incorporating both phrases as the Seq prompt for the sequential approach.
-A.2 Llama 2
-In preliminary experiments, we also tried llama-2-7b-chat and llama-2-13b-chat as LLMs for our methods and found that it is difficult for Llama 2 Touvron et al.
-(2023) to perform topic modeling regardless of prompts and the subset size we use.
-Table 8 shows outputs of Llama 2 when given the Par prompt with the subset size 100 on GoogleNewsT.
-Llama 2 can not make adequate output following our instruction for the number of topics and topic words, while GPT-3.5 and GPT-4 can perform it steadily under the same setting.
-A.3 Subset Size
+A.2 Subset Size
 We assume candidates of the subset size are 250, 500, and 1000.
@@
 We finally select 1000 as the subset size because the performance of each models are relatively high in all metrics under that setting.
+A.3 Llama 2
+In preliminary experiments, we also tried llama-2-7b-chat and llama-2-13b-chat as LLMs for our methods and found that it is difficult for Llama 2 Touvron et al.
+(2023) to perform topic modeling regardless of prompts and the subset size we use.
+Table 8 shows outputs of Llama 2 when given the Par prompt with the subset size 100 on GoogleNewsT.
+Llama 2 can not make adequate output following our instruction for the number of topics and topic words, while GPT-3.5 and GPT-4 can perform it steadily under the same setting.
 Appendix B Experimental Details
-B.1 Implementation Details
+B.1 Preprocessing and Dataset Statistics
+Following Wu et al.
+(2022), the datasets are preprocessed as follows: () characters are converted to lower case; () words of two letters or less are removed; () words appearing less than five times are filtered out.
+Table 9 shows the final statistics of the datasets we use in Section 4 and Section 5.
+B.2 Implementation Details
 We run TSCTM for 200 epochs.
@@
 In the case of data augmentation, we run it with temperature as 0.07, weight contrast as 3.0, and same quant as 0.001.
-For data augmentation, we applied WordNet111111https://github.com/makcedward/nlpaug and Contextual Augmenter33footnotemark: 3 Kobayashi (2018) with 30% word replacement, and filtered low-frequency words as in the preprocessing.
+For data augmentation, we applied WordNet141414https://github.com/makcedward/nlpaug and Contextual Augmenter33footnotemark: 3 Kobayashi (2018) with 30% word replacement, and filtered low-frequency words as in the preprocessing.
 Each Augmenter randomly replaces words in an input text with synonyms defined by WordNet and with words predicted by BERT Devlin et al.
-(2019)121212https://huggingface.co/bert-base-uncased, respectively.
+(2019)151515https://huggingface.co/bert-base-uncased, respectively.
 For gpt-3.5-turbo-0125, gpt-4-0125-preview, and BERTopic, we utilized their original configurations without modification.
-B.2 Examples of Prompts
-Table 10 shows examples of prompts used in the experiment.
+B.3 Examples of Prompts
 Appendix C Result Details
@@
 Following Wu et al. (2022), we randomly select some examples of topics identified by LDA, BERTopic, and our proposed methods with GPT-4.
-C.2 Examples of Topic Words Not Included in the Documents
+C.2 Analysis of Parallel and Sequential Prompting
+Table 2 and Table 3 show that the parallel prompting approach can identify topics with better coherence and document coverage than the sequential prompting one.
+To analyze the superior performance of the parallel approach, we calculate Cv and DC of topics before merging.
+Table 12 shows that Cv and DC scores before merging degrade from those of the parallel approach, demonstrating that the merging process can improve both their coherence and document coverage.
+Additionally, we examine the transition of topics during their processing and then observe that the sequential approach tends to update the previously identified topic very little due to strict adherence to our instruction, leading to lower document coverage compared to the parallel approach.
+Table 13 reports the concrete examples of topics identified for each subset and the final output to demonstrate the processing in our methods.
+In the parallel approach, we find that LLM reasonably merges topics from each subset.
+For instance, bold topics in each subset are merged into one topic in the final output, using words from both subsets.
+In contrast, in the sequential approach, the final output is the same as the topics for the first subset except for only one pair of bold words.
+This indicates that LLMs with the the sequential approach could too strictly keep topics for the previous subset so that they can not output topics that sufficiently reflect the entire set.
+C.3 Examples of Topic Words Not Included in the Documents
 Table 14 shows examples of words not included in the documents outputted in topic modeling on GoogleNewsT.
@@
 However, only a small number of such words were found, and most of them were synonyms, synonyms, derivatives, or related words in the documents.
-C.3 Examples of the Processing
-Table C.3 reports the concrete examples of topics identified for each subset and the final output to demonstrate the processing in our methods.
-In the parallel approach, we find that LLM reasonably merges topics from each subset.
-For instance, bold topics in each subset are merged into one topic in the final output, using words from both subsets.
-On the other hand, in the sequential approach, the final output is the same as the topics for the first subset except for only one pair of bold words.
-This indicates that LLMs with the the sequential approach could too strictly keep topics for the previous subset so that they can not output topics that sufficiently reflect the entire set.
-Appendix D Limitations
-We do not consider it matters whether GPT-3.5 and GPT-4 have seen these datasets in the pre-training; since topic modeling is an unsupervised task, they can not utilize their knowledge about these datasets in our experiment.
-Our approaches perform promisingly in a practical common situation where the number of topics is about 10 but are less reliable with a larger number of topics, such as 50 or 100, due to the difficulty of controlling the LLM’s output format.
+Appendix D Controllability of the Number of Topics
+Unlike conventional topic models that consistently produce the specified number of topics, LLMs may not generate the exact number of topics indicated in the prompt.
+Therefore, to evaluate LLM’s controllability of the number of topics, we investigate its output when the specified number of topics in prompts is set to 5, 10, 15, and 20, while keeping the number of topic words fixed at 5.
+We use the documents set from Section 5 and GPT-4 (gpt-4-turbo-2024-04-09) with Base prompt, running ten trials for each setting.
+The results show that the same number of topics as specified are outputted in all trials, demonstrating GPT-4’s sufficient controllability of the number of topics.
 | ID | Candidates of Base Prompt Template |
@@
 """ [DOCS] """ |
-| Model size | Examples of Llama2 Output |
-| 7B | Topic 1: Top words: relief, challenge, face Topic 2: Top words: welker, concussion, test Topic 3: Top words: live, stream, champion, league Topic 4: Top words: bargain, black, friday, shopping Topic 5: Top words: scotland, independence, white, paper Note: Each topic is represented by 5 top words, which are the most frequently occurring words in the given documents.
-|
-| 13B | Topic 1: Disasters and Relief Efforts Topic 2: Sports and Injuries Topic 3: Technology and Gadgets Topic 4: Politics and Leadership Topic 5: Entertainment and Celebrities |
 | Subset Size | Cv | TU | DC | Fa |
@@
 | 0.562 | 0.984 | 0.197 | 0.984 |
+| Model size | Examples of Llama2 Output |
+| 7B | Topic 1: Top words: relief, challenge, face Topic 2: Top words: welker, concussion, test Topic 3: Top words: live, stream, champion, league Topic 4: Top words: bargain, black, friday, shopping Topic 5: Top words: scotland, independence, white, paper Note: Each topic is represented by 5 top words, which are the most frequently occurring words in the given documents.
+|
+| 13B | Topic 1: Disasters and Relief Efforts Topic 2: Sports and Injuries Topic 3: Technology and Gadgets Topic 4: Politics and Leadership Topic 5: Entertainment and Celebrities |
+| # of | Text | Vocabulary | |
+| Dataset | Documents | Length | Size |
+| Tweet | 2000 | 5.47 | 706 |
+| GoogleNewsT | 11000 | 5.25 | 2376 |
+| StackOverFlow | 19000 | 4.71 | 2544 |
+| 20NG (all samples) | 500 | 147.25 | 2095 |
+| 20NG (computer) | 126 | 98.27 | 1437 |
+| 20NG (politics) | 77 | 296.69 | 1792 |
+| 20NG (science) | 57 | 113.51 | 1672 |
 | ID | Prompt Example |
@@
 | thanksgivukkah hanukkah holiday feast rare |
-| Model | Examples of the topic words not included in the documents |
-| TACTM | twelvemonth sink railway blowout |
-| GPT-3.5 | dorffiefskee broncos patriots health advancement ocean guilty france legal attorney |
-| GPT-3.5 | gree watson advertisement boat funding attorney declared refugees crash digital |
+| Model | Tweet | GoogleNewsT | StackOverFlow | |||||
+| Cv | DC | Cv | DC | Cv | DC | |||
+| GPT-3.5 | 0.532 | 0.366 | 0.517 | 0.569 | 0.464 | 0.634 | ||
+| GPT-4 | 0.580 | 0.395 | 0.523 | 0.665 | 0.519 | 0.747 | ||
 | Subset 1 |
@@
 | christina aguilera national anthem super |
+| Model | Examples of the topic words not included in the documents |
+| TACTM | twelvemonth sink railway blowout |
+| GPT-3.5 | dorffiefskee broncos patriots health advancement ocean guilty france legal attorney |
+| GPT-3.5 | gree watson advertisement boat funding attorney declared refugees crash digital |